12. You have been supplied with the following data for 2 comparable hotel companies (A and B):Company A: Share price: $30; Current EPS: $2; Projected EPS: $3 // Company B: Share Price: $50; Current EPS: $2.5; Projected EPS: $4. Three of the statements below are correct, which ones?

a)      A trades on a greater PE multiple than B

b)      B trades on a greater PE multiple than A

c)      The EPS of A is projected to grow faster than the EPS of B

d)      The EPS of B is projected to grow faster than the EPS of A

e)      A is more expensive than B in terms of valuation

f)       B is more expensive than A in terms of valuation

unanswered
Answer #1

Solution:-

Stock A:

Share price= $30

Current EPS= $2

Projected EPS= $3

EPS growth rate= (3-2)/2= 50%

Existing PE multiple= $30/$2= 15 times

Forward PE multiple= $30/$3= 10 times

Stock B:

Share price= $50

Current EPS= $2.5

Projected EPS= $4

EPS growth rate= (4-2.5)/2.5= 60%

Existing PE multiple= $50/$2.5= 20 times

Forward PE multiple= $50/$4= 12.5 times.

Analysis:

A has an existing multiple of 15 and B has 20. This shows that B trades on a greater PE multiple than A.

B has an EPS growth rate of 60% which is higher than A's EPS growth rate of 50%. This shows that EPS of B is projected to grow faster than that of A.

A has a forward PE multiple of 10 and B has forward multiple of 12.5. This shows that B is more expensive in terms of valuation as compared to A.

On the basis of above, the correct options are (b), (d) and (f).